OpenStudy (anonymous):

2n^2-101=-3 solve the equation by factoring

OpenStudy (anonymous):

7,-7

OpenStudy (anonymous):

n = 7 and -7

OpenStudy (anonymous):

2n^2 - 101 + 3 = -3 + 3, 2n^2 - 98 = 0, 2(n^2 - 49) = 0. n^2 - 49 = 0, (n+7)(n-7)=0. Set both parentheses equal to 0 and solve. n = -7, 7

OpenStudy (anonymous):

you guys are awesome thank you!!